# MMLongBench: Benchmarking Long-Context Vision-Language Models Effectively and Thoroughly

### v1

@@ 4.2 Can Needle-in-a-Haystack Task Reflect LCVLM’s Overall Long-context Ability?
NIAH tasks fail to reflect overall long-context abilities. As shown in Figure 4, none of the NIAH tasks exhibit strong correlation with the broader set of long-context tasks. This suggests that performance on NIAH tasks may not be a reliable indicator of general long-context capabilities. In particular, Visual Haystack (VH) tasks show especially low correlations due to their high difficulty, as discussed above, which results in limited ability to distinguish between models. In MM-NIAH, counting and reasoning tasks show weak correlations with several downstream tasks, with coefficients below 0.8. The retrieval task also shows weak alignment with ICL performance. Interestingly, simpler tasks – like retrieval with a single needle in unrelated essays – tend to correlate better with diverse task categories, which is consistent with our findings in Figure 2. We further examine the differences between text-based and image-based needles in Section D.2.
@@
4.3 Cross-Category Correlations Identify Long-Document VQA as a Reliable Proxy
We perform a cross-category correlation analysis of model performance. We find that different categories do not consistently show strong correlation ($<0.85$) with each other, as shown in Figure 5. Specifically, VRAG and NIAH closely correlate because retrieval is the central capability of both tasks. A further investigation shows that VRAG achieves its highest correlation (of 0.93) with the retrieval task (MM-NIAH-Ret) in Figure 13, reinforcing the shared emphasis on retrieval. Meanwhile, In contrast, summarization and long-document VQA show a high correlation of 0.88, likely due to their shared input format – image-formatted PDF documents. This suggests that image types affect category correlations. In contrast, ICL tasks show relatively weak correlations with other categories. The ICL tasks evaluate models’ ability to induce new classification rules from numerous exemplars, a skill orthogonal to recalling facts in long contexts. This further demonstrates that model developers should consider various long-context skills to draw a more holistic picture of LCVLMs. See Section D.3 for detailed dataset-level correlations and additional category-wise insights.
Long-document VQA as a reliable proxy for long-context capabilities. As shown in Figure 5, long-document VQA achieves the highest average correlation with other categories, indicating that it is more aligned with the broader range of long-context tasks. For example, questions from LongDocURL [17] cover not only simple retrieval but also complex understanding and reasoning. Meanwhile, long-document VQA exhibits the smallest standard deviation, showing that it is also stable and balanced. Taken together, these findings suggest long-document VQA is a more representative and reliable proxy than the commonly adopted NIAH for reflecting overall system performance, allowing model developers to iterate more rapidly without the overhead of full-scale evaluation.
@@ 4.4 Error Analysis
In Figure 6, we show the performance of using another pipeline for DocVQA on MMLongBench-Doc. Here, we convert PDF-formatted documents to plain text with OCR ($\diamond$ w/ OCR) and feed them to LCVLMs. There is no clear winner between the PDF-formatted and OCR-extracted pipelines across all models. While Qwen2.5-VL models perform better with imaged-formatted PDF documents in most cases, Gemma3-27B prefers plain text for shorter input lengths ($\leq$ 32K). Furthermore, we performed a fine-grained analysis by categorizing examples according to their answer sources into two groups: text-pure and vision-needed. As expected, using PDF documents leads to higher scores in vision-needed cases, whereas plain text yields better performance in text-pure cases, especially with longer inputs (64K and 128K). This suggests that OCR capability remains a bottleneck for current LCVLMs when handling long-context inputs. Future work could explore combining both pipelines to further enhance performance. Meanwhile, when using OCR-extracted text, replacing LCVLMs with the corresponding LLMs, Qwen2.5-7B and Qwen2.5-32B ($\diamond$ w/ LLM), yields better results in text-pure cases of DocVQA.
@@ 5 Conclusion
In this work, we have introduced MMLongBench, the first comprehensive benchmark for evaluating long-context vision–language models (LCVLMs) across a wide spectrum of downstream tasks. By covering five distinct task categories—while unifying cross-modal token counting and standardizing context lengths, MMLongBench provides a rigorous, extensible foundation for diagnosing the strengths and weaknesses of frontier LCVLMs. Our evaluation of 46 models reveals that i) evaluation on a single task does not reliably predict overall long-context capability, ii) even frontier models face significant challenges, particularly in OCR accuracy and cross-modal retrieval, and iii) models endowed with enhanced reasoning mechanisms (e.g., “thinking” variants) consistently outperform their base counterparts in long-context settings. Looking forward, we hope MMLongBench will serve as a standard yardstick for the community to benchmark new LCVLMs and to drive research on more efficient vision-language token encodings, more robust position-extrapolation schemes, and improved multi-modal retrieval and reasoning capabilities.
@@ Acknowledgments
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China. We also thank the support from NVIDIA AI Technology Center (NVAITC) and the valuable suggestions on the design of this benchmark provided by Yuxiang Wu, Shizhe Diao, and Hongming Zhang.
@@ A.1 Visual Retrieval-Augmented Generation
InfoSeek [22] is a large-scale dataset for factual knowledge-based VQA featuring long-tail entities from Wikipedia [85]. For InfoSeek, we use the lead section from Wikipedia as gold passages, which is the first section on each page and serves as a summary of the whole page. The lead section may be long, so we chunk it into multiple 100-word passages. We remove all the queries whose corresponding lead section does not contain the correct answer.
ViQuAE [23] repalces the named entities in questions from TriviaQA [72] with corresponding entity images from Wikimedia Commons [^2] https://commons.wikimedia.org/. We obtain gold passages for each question from the KILT benchmark [71], which provides human annotations of gold passages for queries in TriviaQA.
@@
We populate the context with hard negative passages from Wikipedia, and the version we used is the Wikipedia 2019-08-01 dump [71]. We follow the KILT benchmark to preprocess Wikipedia articles into 100-word passages. For retrieval, we adopt a retrieval-and-rerank pipeline, where BM25 is first used for coarse retrieval, followed by reranking with dense embedding from Alibaba-NLP/gte-large-en-v1.5 [86]. Here, we replace the image in each question with its original entity name for better retrieval accuracy, because text-based retrieval achieves higher recall and provides harder distractors.
Previous work [20] shows that this pipeline presents a significantly greater challenge than randomly sampled passages. Also, using a real embedding model for retrieval is consistent with various downstream visual retrieval-agumented generation tasks and can better reflect downstream application performance.
@@ A.2 Needle-in-a-Haystack
For the Visual Haystack dataset [16], we directly use the needles and target objects from the original dataset. Then, we change the number of negative distractors to build long-context examples with a given input length $L$. Note that the original dataset simply reports the image number as context length, ignoring different image sizes. Here, we count image tokens based on patches split by vision encoders as discussed in Section 3.2. There are two tasks in the dataset: VH-Single and VH-Multi, where the target objects are contained in a single-needle image or multiple-needle images, respectively.
@@
We report the accuracy on Visual Haystack, exactly the same as in the original work. The MM-NIAH dataset contains three different tasks: needle retrieval, counting, and reasoning. We use MM-NIAH-Ret, MM-NIAH-Count, and MM-NIAH-Reason as their abbreviations, respectively. In each task, there are both text-needle and image-needle examples. In MM-NIAH-Ret and MM-NIAH-Reason, we use substring exact matching (SubEM) for text-needle examples and accuracy for image-needle examples, exactly following the original paper [18]. In MM-NIAH-Count, we find that the soft accuracy metric proposed in the original paper [18] can be exploited: simply predicting a list of zeros ([0, 0, …]) results in a score over 30 on image-needle examples. Thus, we report the accuracy of the total count of the needle in the haystack instead of comparing the list of needle counts, which we find is more reliable. Last but not least, we sample text-needle and image-needle examples evenly in all three tasks.
@@ A.3 Many-Shot In-Context Learning
We include Stanford Cars [24], Food101 [25], SUN397 [77], and iNat 2021 [27]. At the input length of 128K tokens, we randomly sample 50 different classes from each dataset because we only put about 500 images in the context window. With 50 classes, we can ensure that there are about 10 exemplars from each class, which is sufficient. For iNat 2021, since the dataset contains substantially more classes (over 10,000 species), we randomly sample 50 classes from the “Birds” supercategory and 50 classes from the “Plants” supercategory. For every single example, all the exemplars and the test image are either from the “Birds” classes or the “Plants” classes, ensuring the task remains a 50-way classification problem. Meanwhile, for shorter input lengths, we need to reduce the class number to ensure sufficient shots per class. Specifically, we randomly sample 5, 10, 20, and 40 classes for the input length of 8K, 16K, 32K, and 64K tokens. With those class numbers, we find that the number of exemplars per class is similar to that when there are 128K classes.
@@
We employ a label mapping strategy to ensure that models perform classification based on in-context exemplars instead of relying on their pre-trained knowledge. Each label is randomly mapped to an integer $i\in\{0,1,\ldots,N-1\}$, where $N$ is the number of classes, following established practices [87, 20]. Throughout the evaluation, we provide models with images and their corresponding integer labels. Following Li et al. [75], we arrange exemplars into demonstration rounds, each of which includes exactly one exemplar per label in a random order. We concatenate these demonstration rounds, with the last round truncated, to build examples of input length $L$. Also, the label distribution is balanced in all datasets and input lengths.
@@ A.4 Summarization
Both datasets are constructed using the OCR-extracted plain text as the input. In our evaluation, we replace the OCR-extracted plain text with the original PDF-formatted documents. We screenshot each page of each PDF-formatted document with 144 DPI, following common practices [5]. Different from previous works [5, 17], we do not concatenate images to reduce the token numbers and instead directly feed them into LCVLMs since we are stress-testing the model’s long-context capability.
@@
To control the input length $L$, we truncate document pages from the end. When there are multiple documents in Multi-LexSum, we truncate each document evenly from the end. Additionally, we discard examples that exceed the 128K context length by more than 24K tokens, as adding them would require truncating too many pages to fit within the context window. In this way, we can avoid confounding effects on model performance caused by the loss of key information during page truncation.
@@
Model-Based Metric
The N-gram overlap metrics, such as ROUGE [88], have long been condemned for their poor correlation with human judgment for long-form generation [89, 90]. To ensure reliable evaluation for summarization, we adopt the reference-based LLM evaluation method proposed in HELMET [20]. Specifically, we first break down the gold reference summary into a set of atomic claims with GPT-4o, following prior work [91, 92, 93]. Next, we ask the model to check for three properties of model predictions: precision, recall, and fluency. We utilize GPT-4o to assess if each sentence in the generated summary is supported by the gold reference (precision) and if each claim is present in the generated summary (recall). The F1 score is computed from the recall and precision. We also prompt GPT-4o to assess the fluency of the generated summary. The fluency is assigned a value of 0 if the output is incoherent, incomplete, or repetitive, and a value of 1 if it is fluent and coherent. The final score is the product of fluency and F1 score.
@@
We manually checked 100 atomic claims from 25 Multi-LexSum summaries and another 100 claims from 25 GovReport summaries. We found that only one claim was not factually accurate. Then, we checked the coverage of the claims and found no key facts were missing. This manual verification shows that GPT-4o is virtually always reliable for the decomposition task. For Multi-LexSum, we follow HELMET [20] and use the short summary to obtain atomic claims, where the dataset also provides a long and tiny summary for each case.
@@
From quantitative analysis, we found that GPT-4o can consistently distinguish fluent and non-fluent outputs. The agreement between the model judgements and human judgements is 100% for randomly sampled outputs from GovReport and Multi-LexSum. Then, we sample 10 generated summaries for both GovReport and Multi-LexSum (20 in total) and check 5 atomic claims evaluations for each summary. The models we used are Gemini-2.5-Pro and Qwen2.5-VL-32B. We follow a similar procedure and manually check the precision and recall of those sampled summaries. For precision, we observed Cohen’s $\kappa=0.90$ for GovReport and $\kappa=0.89$ for Multi-LexSum, suggesting almost perfect agreement. Meanwhile, for recall, we observed Cohen’s $\kappa=0.90$ for GovReport and $\kappa=93$ for Multi-LexSum, which are also near perfect.
Inspecting the disagreements, we find that most disagreements come from partially supported cases. We identified two common underlying reasons for partially supported cases when measuring precision and recall, respectively. First, a sentence in a generated summary may contain two points: “While agencies generally documented their review, inconsistencies and documentation gaps existed.” We found the reference summary only supports “inconsistencies and documentation gaps existed,” and the “agencies generally documented” part is an entailment inferred by the model. Such inferred (entailed) information causes a lot of partially supported cases when measuring precision. Second, the claims in the gold reference may include specific details, such as some geographic locations or organization names. These details may not be explicitly mentioned in the generated summary, causing the partially supported cases for recall.

### v2

@@ 4.2 Can Needle-in-a-Haystack Tasks Reflect LCVLM’s Overall Long-Context Ability?
NIAH tasks fail to reflect overall long-context abilities. As shown in Figure 4, none of the NIAH tasks exhibit strong correlation with the broader set of long-context tasks. This suggests that performance on NIAH tasks may not be a reliable indicator of general long-context capabilities. In particular, Visual Haystack (VH) tasks show especially low correlations due to their high difficulty, as discussed above, which results in limited ability to distinguish between models. In MM-NIAH, counting and reasoning tasks show weak correlations with several downstream tasks, with coefficients below 0.8. The retrieval task also shows weak alignment with ICL performance. Interestingly, simpler tasks — like retrieval with a single needle in unrelated essays — tend to correlate better with diverse task categories, which is consistent with our findings in Figure 2. We further examine the differences between text-based and image-based needles in Section D.2.
@@
4.3 Weak Correlation Across Categories Calls for Diverse Evaluation
We perform a cross-category correlation analysis of model performance. We find that different categories do not consistently show strong correlation ($<0.85$) with each other, as shown in Figure 5. Specifically, VRAG and NIAH closely correlate because retrieval is the central capability of both tasks. A further investigation shows that VRAG achieves its highest correlation (of 0.93) with the retrieval task (MM-NIAH-Ret) in Figure 13, reinforcing the shared emphasis on retrieval. Meanwhile, summarization and long-document VQA show a high correlation of 0.88, likely due to their shared input format — image-based PDF documents. This suggests that image types affect category correlations. In contrast, ICL tasks show relatively weak correlations with other categories. The ICL tasks evaluate models’ ability to induce new classification rules from numerous exemplars, a skill orthogonal to recalling facts in long contexts. This further demonstrates that model developers should consider various long-context skills to draw a more holistic picture of LCVLMs. See Section D.3 for detailed dataset-level correlations and additional category-wise insights.
Long-document VQA is a reliable proxy for long-context capabilities. As shown in Figure 5, long-document VQA achieves the highest average correlation with other categories, indicating that it is more aligned with the broader range of long-context tasks. For example, questions from LongDocURL [17] cover not only simple retrieval but also complex understanding and reasoning. Meanwhile, long-document VQA exhibits the smallest standard deviation, showing that it is also stable and balanced. Taken together, these findings suggest long-document VQA is a more representative and reliable proxy than the commonly adopted NIAH for reflecting overall system performance, allowing model developers to iterate more rapidly without the overhead of full-scale evaluation.
@@ 4.4 Error Analysis
In Figure 6, we show the performance of using another pipeline for DocVQA on MMLongBench-Doc. Here, we convert PDF-formatted documents to plain text with OCR ($\diamond$ w/ OCR) and feed them to LCVLMs. There is no clear winner between the PDF-formatted and OCR-extracted pipelines across all models. While Qwen2.5-VL models perform better with PDF-formatted documents in most cases, Gemma3-27B prefers plain text for shorter input lengths ($\leq$ 32K). Furthermore, we perform a fine-grained analysis by categorizing examples according to their answer sources into two groups: text-pure and vision-needed. As expected, using PDF documents leads to higher scores in vision-needed cases, whereas plain text yields better performance in text-pure cases, especially with longer inputs (64K and 128K). This suggests that OCR capability remains a bottleneck for current LCVLMs when handling long-context inputs. Future work could explore combining both pipelines to further enhance performance. Meanwhile, when using OCR-extracted text, replacing LCVLMs with the corresponding LLMs, Qwen2.5-7B and Qwen2.5-32B ($\diamond$ w/ LLM), yields better results in text-pure cases of DocVQA.
@@ 5 Conclusion
In this work, we have introduced MMLongBench, the first comprehensive benchmark for evaluating long-context vision-language models (LCVLMs) across a wide spectrum of downstream tasks. By covering five distinct task categories—while unifying cross-modal token counting and standardizing context lengths, MMLongBench provides a rigorous, extensible foundation for diagnosing the strengths and weaknesses of frontier LCVLMs. Our evaluation of 46 models reveals that i) evaluation on a single task does not reliably predict overall long-context capability, ii) even frontier models face significant challenges, particularly in OCR accuracy and cross-modal retrieval, and iii) models endowed with enhanced reasoning mechanisms (e.g., “thinking” variants) consistently outperform their base counterparts in long-context settings. Looking forward, we hope MMLongBench will serve as a standard yardstick for the community to benchmark new LCVLMs and to drive research on more efficient vision-language token encodings, more robust position-extrapolation schemes, and improved multi-modal retrieval and reasoning capabilities.
@@ Acknowledgments
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China. We also thank the support from NVIDIA AI Technology Center (NVAITC) and the valuable suggestions provided by Yuxiang Wu, Shizhe Diao, and Hongming Zhang on the design of this benchmark.
@@ A.1 Visual Retrieval-Augmented Generation
InfoSeek [22] is a large-scale dataset for factual knowledge-based VQA featuring long-tail entities from Wikipedia [86]. For InfoSeek, we use the lead section of the Wikipedia page for the named entity in the question image as gold passages, which is the first section on each page and serves as a summary of the whole page. The lead section may be long, so we chunk it into multiple 100-word passages. We remove all the queries whose corresponding lead section does not contain the correct answer.
ViQuAE [23] replaces the named entities in questions from TriviaQA [73] with corresponding entity images from Wikimedia Commons [^2] https://commons.wikimedia.org/. We obtain gold passages for each question from the KILT benchmark [72], which provides human annotations of gold passages for queries in TriviaQA.
@@
We populate the context with hard negative passages from Wikipedia, and the version we used is the Wikipedia 2019-08-01 dump [72]. We follow the KILT benchmark to preprocess Wikipedia articles into 100-word passages. For retrieval, we adopt a retrieval-and-rerank pipeline, where BM25 is first used for coarse retrieval, followed by reranking with dense embeddings from Alibaba-NLP/gte-large-en-v1.5 [87]. Here, we replace the image in each question with its original entity name for better retrieval accuracy, because text-based retrieval achieves higher recall and provides harder distractors.
Previous work [20] shows that this pipeline presents a significantly greater challenge than randomly sampled passages. Also, using a real embedding model for retrieval is consistent with various downstream visual retrieval-augmented generation tasks and can better reflect downstream application performance.
@@ A.2 Needle-in-a-Haystack
For the Visual Haystack dataset [16], we directly use the needles and target objects from the original dataset. Then, we change the number of negative distractor images to build long-context examples with a given input length $L$. Note that the original dataset simply reports the image number as context length, ignoring different image sizes. Here, we count image tokens based on patches split by vision encoders as discussed in Section 3.2. There are two tasks in the dataset: VH-Single and VH-Multi, where the target objects are contained in a single needle image or multiple needle images, respectively.
@@
We report the accuracy on Visual Haystack, exactly the same as in the original work. The MM-NIAH dataset contains three different tasks: needle retrieval, counting, and reasoning. We use MM-NIAH-Ret, MM-NIAH-Count, and MM-NIAH-Reason as their abbreviations, respectively. In each task, there are both text-needle and image-needle examples. In MM-NIAH-Ret and MM-NIAH-Reason, we use substring exact match (SubEM) for text-needle examples and accuracy for image-needle examples, exactly following the original paper [18]. In MM-NIAH-Count, we find that the soft accuracy metric proposed in the original paper [18] can be exploited: simply predicting a list of zeros ([0, 0, …]) results in a score over 30 on image-needle counting. Thus, we report the accuracy of the total count of the needle in the haystack instead of comparing the list of needle counts, which we find is more reliable. Last but not least, we sample text-needle and image-needle examples evenly in all three tasks.
@@ A.3 Many-Shot In-Context Learning
We include Stanford Cars [24], Food101 [25], SUN397 [78], and iNat 2021 [27]. Since the 128K context length can accommodate only about 500 images, 50 different classes are randomly sampled from each dataset. With 50 classes, we can ensure that there are about 10 exemplars from each class, which is sufficient. For iNat 2021, since the dataset contains substantially more classes (over 10,000 species), we randomly sample 50 classes from the “Birds” supercategory and 50 classes from the “Plants” supercategory. For every single example, all the exemplars and the test image are either from the “Birds” classes or the “Plants” classes, ensuring the task remains a 50-way classification problem. Meanwhile, for shorter input lengths, we need to reduce the class number to ensure sufficient shots per class. Specifically, we randomly sample 5, 10, 20, and 40 classes for the input length of 8K, 16K, 32K, and 64K tokens. With those class numbers, we find that the number of exemplars per class is similar to that when there are 128K tokens.
@@
We employ a label mapping strategy to ensure that models perform classification based on in-context exemplars instead of relying on their pre-trained knowledge. Each label is randomly mapped to an integer $i\in\{0,1,\ldots,N-1\}$, where $N$ is the number of classes, following established practices [88, 20]. Throughout the evaluation, we provide models with images and their corresponding integer labels. Following Li et al. [76], we arrange exemplars into demonstration rounds, each of which includes exactly one exemplar per label in a random order. We concatenate these demonstration rounds, with the last round truncated if needed, to build examples of input length $L$. Thus, the label distribution is balanced in all datasets and input lengths.
@@ A.4 Summarization
Both datasets are initially constructed using the OCR-extracted plain text as the input. In our evaluation, we replace the OCR-extracted plain text with the original PDF-formatted documents. We screenshot each page of all PDF-formatted documents with 144 DPI, following common practices [5]. Different from previous works [5, 17], we do not concatenate all screenshot pages into one or a few images to reduce the token numbers but instead directly feed them into LCVLMs since we are stress-testing the model’s long-context capability.
@@
To control the input length $L$, we truncate document pages from the end. When there are multiple documents in Multi-LexSum, we truncate each document evenly from the end. Additionally, we discard examples that exceed the 128K context length by more than 24K tokens, as adding them would require truncating too many pages to fit within the context window of 128K. In this way, we can avoid confounding effects on model performance caused by the loss of key information due to severe page truncation.
@@
Model-Based Metric.
The N-gram overlap metrics, such as ROUGE-L [89], have long been condemned for their poor correlation with human judgment for long-form generation [90, 91]. To ensure reliable evaluation for summarization, we adopt the reference-based LLM evaluation method proposed in HELMET [20]. Specifically, we first break down the gold reference summary into a set of atomic claims with GPT-4o, following prior work [92, 93, 94]. Next, we ask the model to check for three properties of model predictions: precision, recall, and fluency. We utilize GPT-4o to assess if each sentence in the generated summary is supported by the gold reference (precision) and if each atomic claim in the gold reference is present in the generated summary (recall). The F1 score is computed from the recall and precision. We also prompt GPT-4o to assess the fluency of the generated summary. The fluency is assigned a value of 0 if the output is incoherent, incomplete, or repetitive, and a value of 1 if it is fluent and coherent. The final score, Fluency-F1, is the product of fluency and F1 score.
@@
We manually checked 100 atomic claims from 25 Multi-LexSum summaries and another 100 atomic claims from 25 GovReport summaries. We found that only one claim was not factually accurate. Then, we checked the coverage of the claims and found no key facts were missing. This manual verification shows that GPT-4o is virtually always reliable for the decomposition task. For Multi-LexSum, we follow HELMET [20] and use the short summary to obtain atomic claims, where the dataset also provides long and tiny summaries for each case.
@@
Quantitatively, we found that GPT-4o can consistently distinguish fluent and non-fluent outputs. The agreement between the model judgements and human judgements is 100% for randomly sampled outputs from GovReport and Multi-LexSum. Then, we sample 10 generated summaries for both GovReport and Multi-LexSum (20 in total) and check 5 atomic claims for each summary. The generated summaries are produced by Gemini-2.5-Pro and Qwen2.5-VL-32B. We follow a similar procedure to the GPT-4o evaluation and manually check the precision and recall of those sampled summaries. For precision, we observed Cohen’s $\kappa=0.90$ for GovReport and $\kappa=0.89$ for Multi-LexSum, suggesting almost perfect agreement. Meanwhile, for recall, we observed Cohen’s $\kappa=0.90$ for GovReport and $\kappa=0.93$ for Multi-LexSum, which are also near perfect.
Qualitatively, inspecting the disagreements, we find that most disagreements come from partially supported cases. We identified two common underlying reasons for partially supported cases when measuring precision and recall, respectively. First, a sentence in a generated summary may contain two points: “While agencies generally documented their review, inconsistencies and documentation gaps existed.” We found the reference summary only supports “inconsistencies and documentation gaps existed,” and the “agencies generally documented” part is an entailment inferred by the model. Such inferred (entailed) information causes a lot of partially supported cases when measuring precision. Second, the claims in the gold reference may include specific details, such as some geographic locations or organization names. These details may not be explicitly mentioned in the generated summary, causing the partially supported cases for recall.